# Explaining Context Length Scaling and Bounds for Language Models

### v2

@@ 4.5 Optimal Context Length and Training Dataset Size: Synthetic Dataset
To validate the idea that longer context lengths are suitable for larger datasets, we conduct experiments on our proposed synthetic dataset with a fixed size MLP. During our experiments, we change the dataset size and context length used, and obtain validation Binary Cross Entropy loss vs. Context Length, as shown in the lower figure in Figure 1. (We also plot $y=-\log(1/2)$ as the loss of random guess in Figure 1.)
@@
5 Related Work
5.1 Enlarging Context Length for LMs
Previous work has made attempts to enlarge the context length of Language Models. Work represented by RoPE(Su et al., 2023) uses rotary positional embedding to support generalizing LMs to longer context in inference compared to the training process. These work uses modified positional embeddings to model the relative position dependency in attention mechanism.
There is also work about enhancing long context understanding and exploring Scaling Laws for context length(Xiong et al., 2024). These work utilize an adjusted pretraining and instruction-finetuning process with more long-context data to enhance the models’ ability on long contexts.
Other work modifying architectures has also been proposed to enhance long context modeling or to simplify deployment of long context LLMs. For example, (Tang et al., 2024) proposes a training-free RazorAttention algorithm to largely compress the KV cache while maintaining performance unchanged.
Architectures and inference methods have been proposed to reduce inference time and memory cost for Language Models, represented by a series of linear transformer or RNN-based methods.(Katharopoulos et al., 2020; Gu & Dao, 2024; Sun et al., 2024). These methods, largely reducing the computational cost and memory usage for long input contexts, have displayed margin ahead of traditional attention-based langauge models for long context inference.
Currently a common practice to train very large Language Models supporting long context is to use pretrain the model with shorter contexts, then finetune them with longer contexts, as presented in tech reports of LLaMa-3(Grattafiori et al., 2024) and DeepSeek-v3(DeepSeek-AI et al., 2024).
5.2 Irrelevant Long Context hurts performance of LMs
Besides context length scaling with relevant contexts, previous researches have studied how LLMs perform for long irrelevant contexts. As an example, (Levy et al., 2024) studies the performance of current LLMs on an adjusted version of ‘needle in a haystack task, where two pieces of key information are embedded into a long text corpora and a question related to both is asked, similar to that presented in Figure 6. The conclusion of these work is that LLMs would perform worse when there is too much irrelevant information.
5.3 Long Context in another field: Time Series Forecasting
Context length, representing the length of input context, is not unique to Nature Language. For time series forecasting, where machine learning plays an important row, there is also work discussing the impact of context length, represented by (Shi et al., 2024). These investigations find that there exists an optimal look-back horizon, which increases with dataset size. However, time series datasets are relatively small compared to NLP datasets, and thus whether this conclusion holds on NLP remains an open problem for this work to study.
5.4 Related Theories for Scaling Laws
Since the discovery of Scaling Laws for Large Language Models (Kaplan et al., 2020) or even earlier, there has been theoretical work trying to explain why model performance could benefit from more data points and more model parameters. For exmaple, (Sharma & Kaplan, 2022) studies the dataset and model scaling from the data manifold perspective.
Specially for Language Models, there is also previous work proposing all kinds of theoretical models. For example, (Michaud et al., 2024) proposes a feature-quant based theory; (Aghajanyan et al., 2020) views the effect of fine-tuning from the intrinsic dimension perspective; (Havrilla & Liao, 2024) proposes to understand scaling with intrinsic dimensions.
@@ 6.1 Conclusion
In Section 2, we propose assumptions related to the relationship between CE Loss, intrinsic dimension and context length. We derive a linear relation between CE loss and Intrinsic Dimension, and study the impact of context length to intrinsic dimension. We further investigate the relationship between intrinsic dimension, context length, and Entropy in intrinsic space in the appendix C from an Entropy perspective.
We also conduct experiments on real data (Section 2, Section 3) and synthetic data (Section 4), on measuring Intrinsic Dimension and on the relationship between Cross Entropy Loss (Bayes Risk and Approximation Loss), Context Length and Intrinsic Dimension.
As a correlation of our theory, for the training-till-overfitting setting, there exists an optimal context length that increases with dataset size in the pretraining process. This is also validated in Section 3 and Section 4.5.
We hope our work may provide insight for future work about long context Language Models, or about Physics for Language Models.
@@ 6.2 Limitations and Future Work
In Section 3 we mainly discuss the case for pretraining with long context, but common practice is to train with shorter contexts and then do long-context-training with longer contexts (DeepSeek-AI et al., 2024). We leave these to future work.
Our theory starting from Intrinsic Dimension only holds with assumptions in Section 2; and in Appendix C we use the perspective of Entropy and possible states number in Intrinsic Space to (partially) explain our assumptions and measurements w.r.t. Intrinsic Dimension. We hope future work may try to propose even more fundamental theories to explain our Intrinsic Dimension measurements and assumptions.
Though not discussed in detail in this paper, there are problems that could be potentially explained by our theoretical framework. For example, in the measurement conducted in this work, the Intrinsic Dimension with increasing context length converges to a finite point, which could potentially explain why RNN models(Gu & Dao, 2024; Peng et al., 2023; Sun et al., 2024) with limited hidden state can be good Language Models.
Statement for Possible Impact
This paper presents work whose goal is to advance the field of Machine Learning (or more specifically, about understanding the impact of long contexts in Language Modeling). There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Definition of Cross-Entropy loss used in this work
It is well-known that
$H_{org}(P,Q)=\sum_{x}$ $-P(x)\log Q(x)$ (8)
@@ Appendix A Definition of Cross-Entropy loss used in this work
where $x_{a:b}$ denotes $x_{a},x_{a+1},\ldots,x_{b-1}$. Although this should be the definition of $H(P,Q)$, it is common practice to calculate perplexity of Language Models with its input as GT lables (e.g. in technical report of LLaMa-3(Grattafiori et al., 2024)), in other words,
$H_{exp}(P,Q)$ (9)
$=$ $\sum_{x}-P(x)\log Q(x)$
@@
Appendix B Detailed Derivation for KL distance in Section 2.2
Based on the definition in Appendix A, we derive $D_{KL}=(dim(\infty)-dim(l))*S$ here.
$D_{KL}(P,P_{l})$ (11)
$=E_{x_{-\infty:0}}\sum_{x_{0}}D_{KL,x_{0}}(P(x_{0}|x_{-\infty,0}),P_{l}(x_{0}|x_{-\infty,0}))$
$=(dim(\infty)-dim(l))*S$
This is because we assume that each dimension in the intrinsic space can store $S$ bits of information.
Another intuitive example is: assuming that the vocab is an integer from 00 to $12^{dim(\infty)*S}-$ 1. assuming $P(x_{0}|x_{-\infty:0})=\delta_{x_{0},y}$, that is, the next token given $x_{-\infty:0}$ is sure to be $y$. $y$. For $P_{l}(x_{0}|x_{-\infty:0})$, the first $dim(l)S$ digits of the integer (in binary representation) are known, but the remaining $(dim(\infty)-dim(l))S$ digits are unknown, making a guess in these numbers yield $P_{l}(x_{0}|x_{-\infty:0})=1/2^{S*(dim(\infty)-dim(l))}$. Thus, $D_{KL,x_{0}}(P(x_{0}|x_{-\infty,0}),P_{l}(x_{0}|x_{-\infty,0}))=1*\log 1/(1/2^{S*(dim(\infty)-dim(l))})=S*(dim(\infty)-dim(l))$.
Appendix C Entropy perspective: hyper-volume in Intrinsic Space
@@
C.1 Bayes Risk from an Entropy-in-Intrinsic-Space perspective
@@
Here we derive similar results as in Section 2.2, but not from an Intrinsic-Dimension perspective, rather we derive it from an Entropy perspective.
@@
  • Assumption 1. Entropy of the Bayes Model $\lim_{l\rightarrow\infty}H_{ntp}(l)=H_{ntp}(\infty)$ is finite, which is the Entropy of next token prediction of language itself.
@@
  • Assumption 2. $\forall l_{1},l_{2}\text{ such that }l_{1}<l_{2}$, $H_{ntp}(l_{1})<H_{ntp}(l_{2})$. This is because a longer context contains more information about the next possible token.
To simplify deduction, we further assume that,
  • Assumption 3. ${D_{KL}}_{,ntp}(P,P_{l})\approx H_{ntp}(P)-H_{ntp}(P_{l})+Const$. This means that the KL divergence between the Bayes LM of infinite context length and the Bayes LM of context length $l$ can be written approximately in the form of entropy difference.
@@
  • Assumption 4. The Entropy w.r.t. Next Token Prediction, i.e., $H_{ntp}(P_{l})$, is linear with the Entropy in the Intrinsic Space of the Bayes Model, i.e., $H_{IS}$.
@@
Entropy of a system can be defined as $H=\log\Omega$ where $\Omega$ is the possible number of states of the system. Similar to the calculation process of Entropy in Statistical Physics(Landau & Lifshitz, 1980), we define Entropy of Intrinsic Space as:
$H_{IS}$
$=$ $\log\Omega$
@@
Here $h$ is the ‘plank constant’, meaning that one state corresponds to a unit hyper-volume of $h^{dim}$ in the Intrinsic Space. A different value of $h$ would only add a constant to $H_{IS}$ and would not affect change in Entropy. Thus, we use $\sum_{i}\log rel\_eigval_{i}$ as Entropy in Intrinsic Space. Moreover, experiments in Figure 10 show that relative eigenvalues increase in the same ratio as the context length increases. This means the number of states in any subspace would scale proportionally with respect to the number of states in the whole space, that is: $H_{subspace}=H_{IS}*dim(subspace)/dim(IS)$, hence Entropy of subspaces are linear with respect to each other.
Though related, Entropy in Intrinsic Space does not equal to Entropy in the next token prediction task. From the probability perspective, let $dec(x)$ be the next decoded token for some point $x$ in the intrinsic space, we have: $H_{IS}=\sum_{x\in IS}-P(x)\log P(x)$, while $H_{ntp}=-\sum_{v\in vocab}P(v)\log P(v)$ where $P(v)=\sum_{x\in IS,dec(x)=v}P(x)$. $H_{ntp}$ is a coarse-grained Entropy compared to $H_{IS}$. $H_{IS}$ contains important information on previous tokens that are important for the prediction of future tokens, while $H_{ntp}$ is related only to the next token.
@@
C.2 Bridging the gap between Intrinsic Dimension explanation and Entropy explanation
@@
Here, starting from previous assumptions and measurements w.r.t. Entropy in Intrinsic Space, we explain why CE is linear w.r.t. Intrinsic Dimension measured in Section 2.2, for $idx>500$. We see in Figure 3 that for $idx>500$, the eigenvalues mainly follow an exponential decay:
$releigval_{ctl,\ idx}=releigval_{ctl,\ 0}*\exp\{-\alpha_{ctl}*idx\}\text{, forcertain context length}$
@@
We also now from the previous result that for different context lengths, the relative eigenvalues increase almost in the same proportion, especially for $idx>1000$. That is, $\alpha_{ctl}=\alpha$, and $releigval_{ctl,\ idx}=\gamma(ctl)*releigval_{\infty,idx}$.
@@
For the subspace for the next token prediction task, the entropy should be proportional to log of volume in the subspace; hence it should be proportional to $m*\log\gamma(ctl)$ where $m$ is the dimension of this exact subspace. That is:
$H_{subspace}(ctl)=m\log\gamma(ctl)+Const$ (12)
@@
For some certain threshold $thres$, we use, for some context length, the measured intrinsic dimension is:
$releigval_{\infty,0}*\gamma*\exp\{-\alpha*dim(ctl,thres)\}=thres,$
@@
hence $\gamma(ctl)=thres/releigval_{\infty,0}*\exp\{\alpha*dim(ctl,thres)\}$. Plugging this into Equation (12) we have:
$CE=-H_{subspace}(ctl)+Const=-m\alpha*dim(ctl,thres)+Const(thres).$ (13)
@@
Thus, we derive our assumptions in Section 2, where $s=m\alpha$. Equation (13) can also be validated in the fourth part of Figure 4, where the measured Intrinsic Dimensions (for $idx\geq 500$) are measured in the exponential decay area, and share similar slopes w.r.t. CE Loss.
C.3 Synthetic dataset: Entropy in Intrinsic Space, and Entropy for output layer

### v3

@@ 4.5 Optimal Context Length and Training Dataset Size: Synthetic Dataset
To validate the idea that longer context lengths are suitable for larger datasets, we conduct experiments on synthetic dataset with fixed size MLP. During experiments, we change the dataset size and context length used, and obtain validation Binary Cross Entropy loss vs. Context Length, as shown in the lower figure in Figure 1. (We also plot $y=-\log(1/2)$ for random guess in Figure 1.)
@@ 5.1 Conclusion
In Section 2, we propose assumptions related to the relationship between CE Loss, intrinsic dimension and context length. We derive a linear relation between CE loss and Intrinsic Dimension, and study the impact of context length to intrinsic dimension. We further investigate the relationship between intrinsic dimension, context length, and Information Entropy in Intrinsic Space in Appendix B from an Entropy perspective. We provide formal definitions of assumptions and derivations of important theorems in Appendix C.
We also conduct experiments with real data (Section 2, Section 3) and synthetic data (Section 4 and Section B), on measuring Intrinsic Dimension and on the relationship between Cross Entropy Loss (Bayes Risk and Approximation Loss), Context Length and Intrinsic Dimension.
As a correlation of our theory, there exists an optimal context length that increases with dataset size in pretraining process. This is validated in Section 3 and Section 4.5. We hope our work may provide insight for future work about long context Language Models, or about Physics for Language Models.
@@ 5.2 Limitations and Future Work
In Section 3 we mainly discuss the case for pretraining with long context, but common practice is to train with shorter contexts and then do long-context-training with longer contexts [19]. We leave these to future work.
Our theory starting from Intrinsic Dimension only holds with assumptions in Section 2; and in Appendix B we use the perspective of Entropy and possible states number for Data Manifold to (partially) explain our assumptions and measurements w.r.t. Intrinsic Dimension. We hope future work may try to propose even more fundamental theories to explain our Intrinsic Dimension measurements.
In our work, similar to several previous work[10, 20], we explain the impact of context length scaling from the perspective of Intrinsic Space (or Data Manifold), which is related not only to data, but also potentially to the neural network (that maps the data into such Intrinsic Space) and the prediction task[10]. Our explanation leans toward how the model represents the data in its intrinsic space and is hence more related to real language models, meanwhile other types of more model-agnostic explanations might also be proposed.
Appendix A Definition and Properties of Cross Entropy loss
A.1 Definition of Cross Entropy Loss discussed in this work
It is well-known that the original definition of Cross Entropy between two sequential distributions $P$ and $Q$: $H_{org}(P,Q)$ should be:
$H_{org}(P,Q)=\sum_{x}$ $-P(x)\log Q(x)$
@@ A.1 Definition of Cross Entropy Loss discussed in this work
where $x_{a:b}$ denotes $x_{a},x_{a+1},\ldots,x_{b-1}$; it is common practice to calculate perplexity of Language Models with its input as GT lables (e.g. in technical report of LLaMa-3[21]), in other words, the experimentally measured Cross Entropy $H_{exp}(P,Q)$ is actually:
$H_{exp}(P,Q)$
@@
A.2 Cross Entropy Loss for Language Model with Context Length $l$
In Equation 8, if $Q_{l}(x_{0}|x_{-\infty:0})$ is a language model with limited context length $l$: $Q_{l}(x_{0}|x_{-\infty:0})=Q_{l}(x_{0}|x_{-l:0})$, we have:
$H(P,Q_{l})$
$=$ $\text{E}_{x_{-\infty:0}}[\sum_{x_{0}}-P(x_{0}|x_{-\infty:0})\log Q_{l}(x_{0}|x_{-l:0})]$
$=$ $-\sum_{x_{-\infty:1}}P(x_{-\infty:1})\log Q_{l}(x_{0}|x_{-l:0})$
$=$ $-\sum_{x_{-\infty:-l}}\sum_{x_{-l:1}}P(x_{-\infty:-l},x_{-l:1})\log Q_{l}(x_{0}|x_{-l:0})$
$=$ $-\sum_{x_{-l:1}}P(x_{-l:1})\log Q_{l}(x_{0}|x_{-l:0})$
$=$ $E_{x_{-l:0}}[\sum_{x_{0}}-P(x_{0}|x_{-l:0})\log Q_{l}(x_{0}|x_{-l:0})]$
$=$ $H(P_{l},Q_{l}).$
Note that $P(x_{0}|x_{-l:0})$ is exactly the Bayes Model with context length $l$. Hence, we have:
$D_{KL}(P,Q_{l})$ $=-H(P)+H(P,Q_{l})$
$=-H(P)+H(P_{l},Q_{l})$
$=-H(P)+H(P_{l})+D_{KL}(P_{l},Q_{l}).$
Specially, if we are calculating the KL Divergence between Nature Language and Bayes Model with context length $l$, thus $Q_{l}=P_{l}$, we have:
$D_{KL}(P,P_{l})=-H(P)+H(P_{l},P_{l})=-H(P)+H(P_{l}).$ (9)
Appendix B Entropy perspective: hyper-volume in Intrinsic Space
@@
B.1 Bayes Risk from an Entropy-in-Intrinsic-Space perspective: Assumptions
@@
Here we derive similar results as in Section 2.2, but not from an Intrinsic-Dimension perspective, rather we derive it from an Information Entropy perspective.
Previously in our main paper and in Appendix A.1, the term ‘Cross Entropy’ refers to difference between two distributions, where larger Cross Entropy loss means worse similarity and hence worse performance. Here we further leverage the concept of ‘Information Entropy’, representing the amount of information carried in the Intrinsic Space: larger information entropy means more information carried. To avoid confusion, we use ‘H’ for ‘Cross Entropy Loss’, and ‘S’ for ‘Information Entropy’.
@@
  • Assumption 1. Information Entropy of Intrinsic Space for Bayes Model $\lim_{l\rightarrow\infty}S(P_{l})=S(P_{\infty})$ is finite, which is the Information Entropy of next token prediction of language itself.
@@
  • Assumption 2. $\forall l_{1},l_{2}\text{ such that }l_{1}<l_{2}$, $S(P_{l_{1}})<S(P_{l_{2}})$. This is because a longer context contains more information.
@@
  • Assumption 3. Linear Entropy Relationship: The Information Entropy w.r.t. Next Token Prediction, defined as $S_{ntp}(P_{l})=H(P_{0})-H(P_{l})$, is linear with the Entropy in the Intrinsic Space of the Bayes Model, i.e., $S_{ntp}(P_{l})=k*S(P_{l})+b$, and $10<k<$ 1. A formal definition can be found in Appendix C.
    $S_{ntp}$ is smaller than $S$ since the Intrinsic Space contains important information on previous tokens that are important for the prediction of future tokens, while $S_{ntp}$ is related only to the next token. The hidden state in RNNs contain more information than only the next token to predict. For example, consider a character-level RNN that predicts the sentence ‘1 + 2 equal_’, the next character to predict is ‘s’, but the hidden state should contain information about answer ‘3’ for the latter tokens.
Please refer to Appendix C for formal definition of Intrinsic Space and formal details.
Assumption 3 assumes a linear relationship between next-token-prediction Cross Entropy Loss and Intrinsic Entropy.
In latter experiments, we experimentally provide an explanation for it. Intuitively, if enlarging the context length would increase each dimension in the Intrinsic Space equally (as shown in Appendix B.2), then the entropy increment of some certain Intrinsic Subspace would be linear to the entropy increment of the Intrinsic Space. If the next-token-prediction task just right corresponds to some certain subspace, then the next-token-prediction entropy would scale linearly with the Intrinsic Space entropy (and entropy for any Intrinsic Subspace as well, as shown in Appendix B.3).
We also show mathematically how to bridge the gap between Intrinsic Dimension explanation and Intrinsic Entropy explanation in Appendix B.4.
In Appendix B.5, we apply this explanation to the synthetic dataset and MLP-based synthetic model we mentioned in our main paper, and in Appendix B.6 we further validate this on a generalized synthetic dataset and Transformer-based synthetic model.
B.2 Experimentally measure Intrinsic Entropy $S$
@@
Entropy of a system can be defined as $S=\log\Omega$ where $\Omega$ is the possible number of states of the system. Similar to the calculation process of Entropy in Statistical Physics[22], we define Entropy of Intrinsic Space as:
$S=$ $\log\Omega$
@@
Here $h$ is the ‘plank constant’, meaning that one state corresponds to a unit hyper-volume of $h^{dim}$ in the Intrinsic Space. A different value of $h$ would only add a constant to $S$ and would not affect change in Entropy. Thus, we use $\sum_{i}\log rel\_eigval_{i}$ as Entropy in Intrinsic Space. Moreover, experiments in Figure 7 show that relative eigenvalues increase in the same ratio as the context length increases. This means the number of states in any subspace would scale proportionally with respect to the number of states in the whole space, that is: $S{subspace}=S*dim(subspace)/dim(intrinsic\ space)$, hence Entropy of subspaces are linear with respect to each other.
B.3 Experimentally measure next-token-prediction Information Entropy $S_{ntp}$
B.3.1 PCA-based Information Entropy Estimation
Though related, Entropy in Intrinsic Space does not equal to Entropy in the next token prediction task. From the probability perspective, let $dec(x)$ be the next decoded token for some point $x$ in the intrinsic space, we have: $S=\sum_{x\in IS}-P(x)\log P(x)$, while $S_{ntp}=-\sum_{v\in vocab}P(v)\log P(v)$ where $P(v)=\sum_{x\in IS,dec(x)=v}P(x)$. $S_{ntp}$ is a coarse-grained Entropy compared to $S$. $S$ contains important information on previous tokens that are important for the prediction of future tokens, while $S_{ntp}$ is related only to the next token.
@@ B.3.1 PCA-based Information Entropy Estimation
B.3.2 Gaussian-KDE based Information Entropy Estimation
In this sub-subsection we use another method for Information Entropy Estimation. As shown in Figure 9, this estimation also aligns well with PCA-based estimation; moreover, such estimated entropy is also linear with respect to Intrinsic Dimension and Cross Entropy Loss.
Figure 9: Gaussian-KDE measured Entropy (10000 samples, auto bandwidth = 0.997756) v.s. PCA-measured Entropy (left), Measrued ID (middle) and Cross Entropy Loss (right).
B.4 Bridging the gap between Intrinsic Dimension explanation and Intrinsic Entropy explanation
@@
Here, starting from previous assumptions and measurements w.r.t. Entropy in Intrinsic Space, we explain why CE is linear w.r.t. Intrinsic Dimension measured in Section 2.2, for $idx>500$. We see in Figure 3 that for $idx>500$, the relative eigenvalues mainly follow an exponential decay:
$releigval_{l,\ idx}=releigval_{l,\ 0}*\exp\{-\alpha_{l}*idx\}\text{, for certain context length}$
where $l$ is the context length, $idx$ is the index of some certain eigen value, and $\alpha_{l}$ is the exponential decay coefficient for this certain context length $l$.
@@
We also see from the previous results (Appendix B.2) that for different context lengths, the relative eigenvalues increase almost in the same proportion, especially for $idx>1000$. That is, $\alpha_{l}\approx\alpha$. We define $\gamma(l)=releigval_{l,0}/releigval_{\infty,0}$ and thus we have: $releigval_{l,idx}=religval_{\infty,0}*\gamma(l)*\exp(-\alpha*idx)$.
@@
For the subspace for the next token prediction task, we denote its dimension to be $m$. Hence, the entropy should be proportional to log of volume in the subspace; that is:
$S_{subspace}(l)$ $=\sum_{idx\in\{\text{dimension of subspace}\}}\log releigval_{\infty,0}\gamma(l)\exp(-\alpha*idx)$ (10)
$=m\log\gamma(l)+Const$
which is the result of the Intrinsic Entropy Explanation.
@@
For Intrinsic Dimension explanation, if we are using some certain threshold $thres$ to measure Intrinsc Dimension, the measured dimension $dim(l,thres)$ should satisfy:
$releigval_{\infty,0}*\gamma(l)*\exp\{-\alpha*dim(l,thres)\}=thres,$
@@
hence the measured dimension is $dim(l,thres)=1/\alpha*(\log\gamma(l)+\log(releigval_{\infty,0}/thres))$. Plugging this into Equation (10) we have:
$L_{CE}=-S_{subspace}(l)+Const=-m\alpha*dim(l,thres)+Const(thres).$ (11)
@@
Thus, we derive our assumptions in Section 2, where $s=m\alpha$. Equation (11) can also be validated in the lower-right part of Figure 3, where the Intrinsic Dimensions (for $idx\geq 500$) are measured in the exponential decay area, and these lines, though measured with different threshold ($thres$), share similar slopes w.r.t. CE Loss (that is not related with threshold, as shown in Equation 11).
B.5 Synthetic dataset: Entropy in Intrinsic Space, and Entropy for output layer